OpenStudy (anonymous):

one dimension of a rectangular is 8 inches and the perimeter is 40 inches. what is the other dimension of the map ?

OpenStudy (doc.brown):

Trace the rectangle with your finger. First you move along the 8inch side, then along the side of unknown length, then along the other 8inch, then another unknown, I called the unknown length "x".

OpenStudy (doc.brown):



OpenStudy (doc.brown):

\[8+x+8+x=40\]Group like terms\[x+x+8+8=40\]simplify\[2x+16=40\]Try to get x by itself by first subtracting 16 from both sides